OpenStudy (anonymous):

x + (-15)= -13 + (-2)

OpenStudy (australopithecus):

Add 15 to both sides of the equation

OpenStudy (australopithecus):

you want to isolate x on one side of the equation

OpenStudy (australopithecus):

remember, +(+1) = +1 -(+1) = -1 +(-1) = -1 -(-1) = +1 or +*+ = + +*- = - -*- = +

OpenStudy (skullpatrol):

Use the idea of balance @studyislandhelp

OpenStudy (skullpatrol):

Whatever you do to one side of an equation you must do the same thing to the other, to keep the equation "balanced".

OpenStudy (solomonzelman):

congrats on the 90, you're green now, @skullpatrol !

OpenStudy (skullpatrol):

Thanks @SolomonZelman :D

OpenStudy (solomonzelman):

yw